OpenStudy (sloppycanada):

Can anyone walk me through solving one of these types of problems? Find the sum of the following infinite geometric series, if it exists. 1.02 + 2.04 + 4.08 + 8.16 +…

OpenStudy (anonymous):

well, it looks like y = x(1.02)

OpenStudy (anonymous):

And the formula for infinite geometric series is S = a1/ 1 - 1/2

OpenStudy (anonymous):

1- r my bad

OpenStudy (kropot72):

The common ratio of the series in the question is r = 2.04/1.02 = 2. The series can be summed to infinity only if r < 1.

OpenStudy (danjs):

The thing blows up